A supermarket organization buys in a particular foodstuff from four suppliers A B,C and D and subjects samples of this to regular tasting tests by expert panels.various characteristics are scored and the total score for the product is recorded. Four tasters a,b,c and d at four sessions obtained the results below. Analayse and comment on these.
Taster a
Session 1 : A 21
Session 2: B 20
Session 3: C 20
Session 4 : D 32

Taster b
Session 1: B 17
Session 2 : D 22
Session 3: A 24
Session 4: C 21

Taster c
Session 1: C 18
Session 2: A 23
Session 3: D 22
Session 4: B 22

Tester d
Session 1: D 20
Session 2: C 19
Session 3: B 19
Session 4: A 26
1
Expert's answer
2020-05-01T17:56:37-0400

taster a, considers product D. To be the best

taster b, considers product A. To be the best

taster c, considers product A. To be the best

taster d, considers product A. To be the best.

According to most, the best product is A

We can calculate the average product rating and the deviation from each of the tasters for the better.

It will be more accurate than just the sum of the readings of different tasters, because The weight of the assessment is different.

For Expert A, the average rating is 23.25. (Product: 4. Summ rating tast is 21+20+20+32=93. 93/4=23,25 - the median rating the product for expert A)

and for expert A, to a better estimate, the deviation is 8.75 (32-23,25)


what about the average estimate +0.344 (product D) (8,75/23,25; deviation/average rating)

For product A: average estimate -0.097 (deviation is -2,25=21-23,25; rating minus average rating; -2,25/23,25=-0.097)

Similar values are for experts: for B - average rating 21 (17+22+24+21=84;/4=21), the deviation is 3 (forbproduct A: 24-21=3), estimate 0.125.

For C -av.r=21,25 (18+23+22+22=85;/4=21,25), dev=1,75 (23-21,25=1,75), est=0.082 (1,75/21,25)

for D av,r=21(20+19+19+26=84:/4=21),dev=5 (26-21=5)est=0.238 (5/21=0.238)

Product A is more successful at 0.348 (summ estimate; -0.097+0.125+0.082+0.238=0.348)./4(number of tast,or number of expert) =0.087